OpenStudy (anonymous):

Determine whether the graph of y = x^2 - 6x + 3 has a maximum or minimum point, then find the maximum or minimum value. What are the x-intercept(s) of the graph of y - 3 = x^2 + 4x? What are the solutions of 5x^2 = -15x? Where does the graph of y = 2x^2 + 7x + 3 cross the x-axis? x^2 - 3x = 28 solve each

OpenStudy (anonymous):

1) y'= 2x-6=0 => x=3 and plug it into y and you get y=9-18+3=-6 (its minimum value) 2) y=0 => x^2+4x+3=0 => x=-3,-1 3)5x^2+15x=0 => 5x(x+3)=0 => x=0,-3 4)y=0 => 2x^2+7x+3=0 => x=-3/2,-1 5) x^2-3x-28=0 => x=-4,7